How many tosses of a fair coin are needed so that the probability of getting atleast one head is 0.875?


1
Expert's answer
2020-07-09T19:26:38-0400

Solution.

Let "p = 0.875" be the probability of getting at least one head. Let "q" be the probability of getting all tails. It is logical that "p + q = 1." The coin is fair, so after tossing "n" times the probability of of getting all tails is "q = 0.5^n". So we get the equation:


"0.875+0.5^n = 1 \quad \Longrightarrow \quad n = \log_{0.5}(1-0.875) = 3"


Answer: 3